the residents of a city voting on whether to raise property taxes the ratio of yes to no votes 3to2 if there were 1608 yes votes what is the total number of votes?
Answer by lwsshak3(11628)

the residents of a city voting on whether to raise property taxes the ratio of yes to no votes 3to2 if there were 1608 yes votes what is the total number of votes?
***
let x=total number of votes
x-1608=number of no votes
1608/(x-1608)=3/2
3216=3x-4824
3x=8040
x=2680
total number of votes=2680